LEP EXPERIMENTS COMMITTEE
Minutes of the 48th meeting,
31 March 1998

OPEN SESSION

1. Summary of the 1998 Chamonix Workshop and
Impact of LHC civil engineering on LEP

J. Poole reported on the conclusions of the Chamonix VIII workshop, January 1998. During the shutdown another 8 sc rf modules
(32 cavities) were installed at LEP and more of the warm copper cavities removed. This brings the total number of sc cavities to 272
and allows the operating energy for 1998 to be raised to ÷s = 189 GeV, allowing for two klystrons to trip without losing the beam. Based on a short but encouraging test in 1997, the machine will be operated with 102/90 optics in 1998. The performance goal is an integrated luminosity of 150 pb-1 per experiment by the end of the year, implying an average luminosity of 1.2 pb-1 per day. There will be four bunches per beam, with a current of about 800 mA per bunch. For Z calibration data, the machine will be operated with either 102/90 or 90/90 optics and eight bunches per beam. To help achieve the luminosity goal for 1998, numerous improvements have been made to raise the operational efficiency of LEP. Among these are modifications to transitions in the vacuum chamber in order to reduce the heat loads from synchrotron radiation.

Concerning the LHC civil engineering, the impact on LEP operation in 1998 and 2000 should be relatively small. However, during 1999 vertical movements of the LEP tunnel by up to 30 mm are expected near the experimental pits for ATLAS and CMS. The largest movements have been concentrated during shutdown periods, but the schedule is subject to change. Frequent vertical realignment of the affected parts of the machine will be necessary during the course of the year, causing an estimated SPS downtime of 36-72 hours and a LEP downtime of 100-125 hours.

2. LEP2 physics jamboree

Each experiment reported samples of about 700 WW events in 1997, which provide a statistical precision on the W mass of about
140 MeV per experiment. The lower mass limits on the SM Higgs are 84-88 GeV per experiment. The lower mass limits on the
lightest Higgs of minimal SUSY (h,A) are 71-76 GeV per experiment for large values of tan b; the low tan b range (0.6-2) is excluded for no-mixing scenarios.

ALEPH reported a tracking mis-alignment which had varied over a period of about 3 months in the second half of the 1997 run. The exact cause is not known. As a result ALEPH requested an additional 0.5 pb-1 Z calibration data near the end of the 1998 run, with the option to request an additional 0.5 pb-1 in the middle of the run if necessary.

3. LEP working group reports

The combined W mass measurement from LEP, 80.35±0.09 GeV, now matches the precision of the Tevatron value, 80.40±0.09 GeV. The effective weak mixing angle measured at LEP is 0.23185±0.00026, to be compared with the SLD value of 0.23084±0.00035.
The discrepancy between the measurements at the two machines has reduced from 2.8 to 2.3 s over the last year. The predicted mass
of the standard model Higgs using all electroweak data (updated since the open session) is mH = (66+75-39) GeV, or mH < 215 GeV
(95% CL).

Concerning the LEP2 energy calibration, a new, independent measurement based on a beam spectrometer is being prepared, aiming
at a precision s = 10 MeV at a beam energy of 100 GeV. This involves measuring the beam trajectory with high-precision beam orbit monitors (sx = 1 mm) either side of a precisely-mapped dipole magnet (DB/B = 10-4). A test installation comprising two concrete magnets and one precision triple BOM assembly has been installed in LEP for studies during 1998. The definitive installation of a precision-mapped steel magnet and a second triple BOM assembly will be made in the 98/99 shutdown.

CLOSED SESSION

Present: J. Drees, L. Foà, G. Goggi, K. Hübner, P.O. Hulth, G. Kantardjian, J. Kirkby (Secretary), K.-H. Kissler,
                W. Lohmann, T. Lohse, M. Mangano, J. May, T. Müller, S. Myers, J. Panman, L. Pape, M. Pohl (part time),
                S. Pokorski, Y. Sirois, P. Sphicas, I. Videau, P. Wells and P. Zerwas (Chairman).

Apologies: S. Bethke, R. Cashmore, J. Dainton, F. Gasparini, R. Marshall.

1. Approval of the minutes of the 47th meeting

The minutes of the 47th meeting (LEPC 97-12/LEPC 47) were approved without modification.

2. Chairman's report

The Research Director was pleased to announce that Roger Cashmore has accepted to be the next Chairman of the LEPC after
Peter Zerwas, whose term ends with the present meeting. He joined with the Chairman in thanking Pippa Wells for her excellent
work as LEP Coordinator, and in welcoming Paris Sphicas as the new LEP Coordinator.

The Chairman summarised the outcome of discussions that had taken place regarding guidelines for the LEP working groups. An initial meeting of the working group chairmen had led to a useful exchange of views; the matter had been referred to the LEP spokesmen for the final decisions. At their subsequent meeting, the spokesmen had decided that drawing up formal guidelines would be unnecessarily rigid, and that regular feedback between the collaborations and their representatives in the working groups was sufficient to resolve any issues that may arise.

The Research Director explained the status of the discussions concerning special contributions from the member- and non-member states for operating LEP in 2000. Following the initial requests made by the DG and the subsequent offers of support, formal commitments have now been requested. It is expected that the exact financial plan for LEP operation in 2000 will be understood before the
June Council meeting. Although the final decision will only be known at that time, the indications are very strong that LEP will
operate in 2000.

3. Discussion on the LEP machine reports

The committee supports the plan that LEP operation in 1998 be aimed towards achieving the highest possible integrated luminosity, even where this would require a small reduction in the peak energy. In this regard, the committee expressed its appreciation of the studies that were reported on detailed improvements in the operational efficiencies of many areas of the machine. Concerning the impact of LHC civil engineering on LEP operation, the committee was pleased to note that both this year and 2000 should be relatively "safe" years. Moreover, the small movements expected in 1998 (restricted to the SPS) will allow an experimental cross-check of the predicted values, and provide useful operational experience for the accelerators.

4. Discussion on the reports from the LEP experiments and working groups

The committee congratulated the combined LEP machine and experimental community for measuring MW to a precision that now matches the Tevatron experiments. The most interesting physics result is the falling predicted mass of the standard model Higgs; using all electroweak data, mH = (66+75-39) GeV, or mH < 215 GeV (95% CL). This raises the prospects for the Higgs existing within the discovery mass limit of LEP2, which is about 105 GeV.

Concerning the LEP2 beam energy calibration, the committee was pleased to hear the progress towards the beam spectrometer, which will provide a measurement of the LEP2 beam energy with a precision goal of s = 10 MeV. This is a factor three better than the present error and probably a factor two better than the ultimate precision of the present technique. Based on measurements with the hardware presently being installed at LEP, there should be a clear understanding by the end of the year whether this precision goal for the beam spectrometer can be achieved.

There was an extended discussion on the LEP data archive: the LEDA project. Whereas all experiments consider an archive is valuable and worth the effort, there were varied views concerning the possible eventual physics use of the archive. Until the latter is more clearly defined, it is difficult to decide on exactly what needs to be archived and how much work and resources are involved. In summary, the LEPC supports in principle the archiving of LEP data. It suggests the working group could extend its membership to include input from a theorist and a non-LEP experimentalist and seek to define real physics goals for the archive, in consultation with the LEP collaborations. This would allow more concrete proposals to be made on what should be archived and how to do it. The Chairman proposed that a meeting be organised in about 2 months time to arrive at a definitive decision how to proceed. The meeting would include the archive working group Chairman, the LEP spokesmen, the Research Director, the IT Division Leader and LEPC representatives.

5. Report from the LEP Coordinator

The LEP Coordinator reported that LEP will operate for 169 days in 1998, comprising 125 d for high energy data, 10 d for Z calibration (assuming 2.5 pb-1 and either the 102/90 or 90/90 optics), 22 d for LEP MD and 12 d for non-LEP stops and MD. The LEP energy calibration runs are included in the LEP MD periods. Setup with beams is expected on 11 May (with the possibility of beam as early
as 4 May). The cycle presently ends on 1 November, with one extra week (45) in reserve as a possible additional week for
LEP MD. A one-week (39) interruption is foreseen when the SPS switches from protons to Pb.

6. ALEPH request for additional Z calibration data

Concerning the request from ALEPH for an additional 0.5 pb-1 Z calibration data near the end of the 1998 run, the other three experiments expressed their support. The LEPC therefore recommends approval of an additional 0.5 pb-1 Z data near the end of the run, subject to the following conditions. Any requests for additional Z calibration data in the middle of the run would require further discussion. The same operational conditions are applied to this Z calibration sample (2 days) as for the 2.5 pb-1 sample (10 days)
at the start of the run; namely if, for unforeseen reasons, it is found to require a significantly longer time then the decision to switch to high-energy running will be taken by the Research Director, in consultation with the LEP spokesmen. The committee also asked that ALEPH make a status report on their alignment problem at the September LEPC meeting and confirm that the additional Z calibration data is still necessary.

7. Discussion on the L3 cosmic ray experiment

The Chairman invited M. Pohl of the L3+C collaboration to make a brief presentation of the L3 cosmic ray experiment, which involves covering the outside of the L3 magnet octants with (mainly recuperated) plastic scintillation counters. The committee expressed its appreciation of the great physics interest of this experiment, which should make a precision measurement of the cosmic muon spectrum from 20 GeV to 2 TeV as a function of momentum, zenith angle and charge. The expected precision is at the per cent level, to be compared with the present systematic uncertainties of 20% or more. It was noted that L3+C is fully compatible with normal L3 operation and that the experiment would terminate at the end of LEP operation, whether it be 1999 or 2000. The committee also understands that the L3+C collaboration undertakes to make no dedicated access requests for their detector during LEP operation. In conclusion, the committee is satisfied that the L3+C experiment can operate parasitically during L3 operation at LEP.

8. Any other business

The Secretary announced that Monique Budel has set up a web site where useful LEPC documents will be made available (http://www.cern.ch/Committees/LEPC/WelcomeLEPC.html).

At the close of the meeting, the Research Director expressed his warmest thanks to the outgoing Chairman, Peter Zerwas. He highlighted the splendid way that Peter had transmitted to the SPC and the Research Board the enthusiasm and physics importance for operating
LEP in 2000. The Research Director expressed his appreciation on behalf of the entire LEP community for the considerable role Peter has played in making LEP operation in 2000 the top priority of CERN's near-term programme.

9. Next LEPC meeting

The dates of the next meeting of the LEPC are Tuesday-Wednesday, 15 - 16 September. The dates of the final meeting in 1998 are 12 - 13 November.
 

DOCUMENTS RECEIVED

[1] L3+C Collaboration, The L3 Cosmic Ray Experiment, LEPC 98-3/ M 116.